OpenStudy (woznichak):

Let f(x) = x − 3 and g(x) = x + 11. Find f(x) ⋅ g(x). Please help!!

OpenStudy (jango_in_dtown):

f(x).g(x)=(x-3)(x+11)= ??

OpenStudy (woznichak):

i dont have any clue.

OpenStudy (jango_in_dtown):

=x(x+11)-3(x+11). now expand

OpenStudy (jiteshmeghwal9):

use distributive property a(b+c)=ab+ac

OpenStudy (jiteshmeghwal9):

to expand *

OpenStudy (woznichak):

so would it is x^2+11-3x+11